OpenStudy (anonymous):

need to be show the steps cuz i keep getting the wrong answer 5[2-(2x-4)]=2(5-3x)

OpenStudy (anonymous):

distrubute 10-10x-20=10-6x -4x=20 x=20/-4 x=-5

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yeah you probably got tripped up distributing the 5. 5*2-5(2x-4), so you gotta distribute it twice